# Limitations of Data-Driven Spectral Reconstruction: Optics-Aware Analysis and Mitigation

### v3

@@ Abstract
Hyperspectral imaging empowers machine vision systems with the distinct capability of identifying materials through recording their spectral signatures. Recent efforts in data-driven spectral reconstruction aim at extracting spectral information from RGB images captured by cost-effective RGB cameras, instead of dedicated hardware.
In this paper we systematically analyze the performance of such methods, evaluating both the practical limitations with respect to current datasets and overfitting, as well as fundamental limitations with respect to the nature of the information encoded in the RGB images, and the dependency of this information on the optical system of the camera.
We find that, the current models are not robust under slight variations, e.g., in noise level or compression of the RGB file. Without modeling underrepresented spectral content, existing datasets and the models trained on them are limited in their ability to cope with challenging metameric colors. To mitigate this issue, we propose to exploit the combination of metameric data augmentation and optical lens aberrations to improve the encoding of the metameric information into the RGB image, which paves the road towards higher performing spectral imaging and reconstruction approaches.
Index Terms:
Hyperspectral imaging, Spectral reconstruction from RGB, Metamerism, Overfitting, Aberration.
@@ I Introduction
Refer to caption fake_and_real_food_ms spectra
@@ II-A Hyperspectral cameras
Conventional hyperspectral imaging systems require filter wheels, liquid-crystal tunable filters, or mechanical motion (e.g., pushbroom) [44, 32] to scan the 3D hyperspectral datacube. To enable snapshot acquisition, coded-aperture snapshot spectral imager (CASSI) [62, 19] has been proposed to achieve high spectral accuracy using spectrum-dependent coded patterns. Recent methods also exploit spectrally encoded point spread functions (PSFs) to computationally reconstruct a hyperspectral image [36, 9, 15]. In general, great efforts have been made to simplify hyperspectral camera hardware by software reconstruction.
@@ II-B Spectral reconstruction from RGB images
A recent trend to solve the snapshot hyperspectral imaging problem is to exploit hyperspectral data with deep neural networks to reconstruct spectral information from RGB images [4]. Owing to the wide availability of RGB cameras, this approach seems to be a promising candidate for hyperspectral imaging if successful. A large number of neural network architectures have been proposed in the past three NTIRE spectral recovery challenges  [5, 6, 7]. Our analysis in this paper focuses on this class of methods to gain insights on their strengths and limitations.
II-C Dataset bias and data augmentation
@@
II-D Metamerism
Metamerism is a physical phenomenon where distinct spectra produce the same color [2] as the high-dimensional spectral space is projected down to three dimensions of a trichromatic vision system (either the human eye or an RGB camera). This phenomenon has been studied in color science [56, 28], spectral rendering [35, 63, 58], and hyperspectral imaging [29, 27]. In hyperspectral imaging [21, 47], it is crucial for many applications to distinguish between metamers or near-metamers (i.e., different spectra that project to similar RGB values). Indeed spectral imaging is usually employed when the RGB color differences between two materials or features are too small to reliably distinguish between them. Therefore, hyperspectral imaging systems require special attention in the system design to acquire accurate spectral signatures [30, 37].
@@ III-A Spectral image formation
In the past NTIRE challenges [5, 6, 7], optical aberrations haven’t been included in the image simulation pipeline. However, the optical system of the RGB camera inevitably introduces spectrally-varying blurs to the spectral images, which is modeled as PSFs. This optical process can be described by a linear matrix-vector product in each spectral band followed by a sum over the spectral dimension. The spectral images through the optical system are ${\mathbf{W}}=\mathrm{diag}\left({\mathbf{A}}{\mathbf{X}}\right)$, where ${\mathbf{A}}\in\mathbb{R}^{KMN\times MN}$ is a block matrix that stacks the spectral PSF matrices vertically, and $\mathrm{diag}(\cdot)$ extracts the diagonal blocks. The final RGB image is then
@@ III-B Hyperspectral datasets and data diversity
Compared to very large color (RGB) image datasets (e.g., ImageNet [23], DIV2K [1]), hyperspectral datasets are far smaller in size, primarily limited by the unavailability of high-quality hyperspectral cameras and the difficulty in acquiring outside the lab with moving target scenes. The largest dataset so far is ARAD1K used in the NTIRE 2022 challenge [7]. In addition, we also include the CAVE [67], ICVL [4], and KAUST [42] datasets that share the same spectral range to extend our experiments. The datasets are summarized in Table I. Although other datasets [16, 18, 46] exist, they cover slightly different spectral bands, making direct comparisons difficult. We therefore restrict our analysis to the datasets listed in the table.
@@ III-C Modeling metamerism
A spectrum ${\mathbf{S}}$ can be projected onto two orthogonal subspaces, one for the fundamental metamer ${\mathbf{S}}^{*}$, and the other for metameric black ${\mathbf{B}}$ [20, 61]. A new metameric spectrum ${\mathbf{S}}^{\prime}$ is then
@@
where ${\mathbf{S}}^{*}={\mathbf{Q}}\left({\mathbf{Q}}^{T}{\mathbf{Q}}\right)^{-1}{\mathbf{Q}}^{T}{\mathbf{S}}$ and ${\mathbf{B}}={\mathbf{S}}-{\mathbf{S}}^{*}$. Since adding metameric black does not alter the RGB color, we can vary the coefficient $\alpha$ to generate different spectra that are all metamers. To avoid negative spectral radiance, we clip the negative values in the generated spectra and re-calculate the RGB colors for the affected pixels. See Appendix for the analysis on the effects of clipping to non-negative values while generating metamer data.
@@ III-D Performance evaluation metrics
${\mathrm{\mathbf{MRAE}}}=\frac{1}{KMN}\sum_{k,i,j}\left|\hat{{\mathbf{X}}}_{k,i,j}-{\mathbf{X}}_{k,i,j}\right|/{\mathbf{X}}_{k,i,j},\$ (4)
@@ III-E Training details
In our study, we conduct all the experiments across various datasets and network architectures. Following the methodologies proposed by the latest champion network MST++ [14], we employ their patch-wise training approach (patches of 128$\times$128 pixels).
@@ III-E3 Training and Validation Procedure
Following the training strategies in MST++ [14], we sub-sample the hyperspectral datacubes and the corresponding RGB images into overlapping patches of 128$\times$128. Spatial augmentations, such as random rotation, vertical flipping, and horizontal flipping are randomly applied to the training patches. In the validation step, we calculate the evaluation metrics (MRAE, RMSE, PSNR, and SAM) on the full spatial resolution for the ARAD1K (482$\times$512), CAVE (512$\times$512), and KAUST (512$\times$512) datasets. Note that this is different from MST++ [14] where only the central 256$\times$256 regions are evaluated. The ICVL dataset has a very large spatial resolution (1300$\times$1392). To be consistent with other datasets, we evaluate ICVL only in the central 512$\times$512 regions.
@@ IV Finding 1: Atypical Overfitting
Although it is well-known that deep neural networks may suffer from overfitting problems, we find that the overfitting behavior in spectral reconstruction is atypical and difficult to notice with standard evaluations. Here we introduce minimalist changes to the ARAD1K dataset used in NTIRE 2022 challenge [7] in three experiments to demonstrate it. We evaluate 11 top-performant neural network architectures, namely MST++ [14], MST-L [13], MPRNet [70], Restormer [68], MIRNet [69], HINet [17], HDNet [31], AWAN [41], EDSR [43], HRNet [72], and HSCNN+ [54].
@@ IV-A Training with less data
First, we make a simple change to the training of the participating networks in the NTIRE 2022 challenge [7]. While keeping all the training settings intact, we randomly choose only 50% or 20% of the original training data, respectively, to train the candidate networks and validate the performance on the original validation data. We illustrate the validation curves for MST++ [14] in Fig. 3. See the Supplemental Material for the results of other networks. We summarize the results for 100% and 50% training data in Table II for all networks.
@@ IV-B Validation with unseen data
In our experiments, we generate new RGB images using the same methodology and calibration data, but different noise and compression settings. Specifically, we use the SRF data for a Basler ace 2 camera (model A2a5320-23ucBAS) known to the networks, and simulate Poisson noise at varying noise levels by controlling the number of photon electrons (npe). We adopt the same rudimentary in-camera image signal processing pipeline. As an illustrative example, we use MST++ [14] in Table LABEL:tab:unseen, Row 1 as a reference for comparison; results for other networks can be found in the Supplemental Material.
First, as a baseline, we consider a noiseless (npe = 0) and aberration-free case with moderate JPEG compression quality (Q = 65), shown in Table LABEL:tab:unseen, Row 2. The results show significant drops in all the performance metrics. Note that the only differences here compared to the challenge dataset are the noise level and compression quality – the base images are identical! This indicates that the network overfits both the noise and JPEG compression parameters.
@@ IV-C Cross-dataset validation
In addition, we inspect the effects of different datasets (cf. Table I) on the performance. We train the MST++ network on the four datasets with the same image simulation parameters. To eliminate the impact of other factors, we choose the ideal noiseless and aberration-free condition without compression. In the validation, we use our trained model on ARAD1K dataset to validate on the other three datasets, respectively. In Table IV, we compare the performance with the models both trained and validated on the original datasets. Results for other networks can be found in the Supplement. They all illustrate the same difficulties in generalization.
TABLE IV: Cross-dataset validation example using MST++ [14].
@@
Even though the imaging conditions are the same and ideal, the network trained on one dataset experiences significant performance drop in all metrics when validated on other datasets. This indicates that the contents of the datasets, as well as the acquisition devices used to capture the datasets, play important roles.

### v4

# Limitations of Data-Driven Spectral Reconstruction: An Optics-Aware Analysis

@@ Abstract
Hyperspectral imaging empowers machine vision systems with the distinct capability of identifying materials through recording their spectral signatures. Recent efforts in data-driven spectral reconstruction aim at extracting spectral information from RGB images captured by cost-effective RGB cameras, instead of dedicated hardware. Published work reports exceedingly high numerical scores for this reconstruction task, yet real-world performance lags substantially behind.
In this paper we systematically analyze the performance of such methods with three groups of dedicated experiments. First, we evaluate the practical overfitting limitations with respect to current datasets by training the networks with less data, validating the trained models with unseen yet slightly modified data, and cross-dataset validation. Second, we reveal fundamental limitations in the ability of RGB to spectral methods to deal with metameric or near-metameric conditions, which have so far gone largely unnoticed due to the insufficiencies of existing datasets. We achieve this by validating the trained models with metamer data generated by metameric black theory and re-training the networks with various forms of metamers. This methodology can also be used for data augmentation as a partial mitigation of the dataset issues, although the RGB to spectral inverse problem remains fundamentally ill-posed.
Finally, we analyze the potential for modifying the problem setting to achieve better performance by exploiting some form of optical encoding provided by either incidental optical aberrations or some form of deliberate optical design. Our experiments show that such approaches do indeed provide improved results under certain circumstances, however their overall performance is limited by the same dataset issues as in the plain RGB to spectral scenario. We therefore conclude that future progress on snapshot spectral imaging will heavily depend on the generation of improved datasets which can then be used to design effective optical encoding strategies. Code can be found at https://github.com/vccimaging/OpticsAwareHSI-Analysis.
@@ II-A Hyperspectral cameras
Conventional hyperspectral imaging systems require filter wheels, liquid-crystal tunable filters, or mechanical motion (e.g., pushbroom) [56, 43] to scan the 3D hyperspectral datacube. To enable snapshot acquisition, coded-aperture snapshot spectral imager (CASSI) [79, 25] has been proposed to achieve high spectral accuracy using spectrum-dependent coded patterns. Based on this hardware architecture, supervised learning (such as TSA-Net [59], BiSRNet [16], VmambaSCI [103], SpeCAT [95]) and unsupervised learning (such as LRSDN [23], SAH-SCI [100], CEINR [58]) algorithms have been proposed in recent years to address the inverse problem on hyperspectral datasets. Other variants, such as dual-camera CASSI [82] and reconstruction algorithms have also been proposed. Another category of emerging methods also exploit spectrally encoded point spread functions (PSFs) to computationally reconstruct a hyperspectral image [47, 10, 20]. Various DOE designs for optimal spectral PSFs, such as equalization DOE [92], non-serial quantization-aware deep optics [81], tunable phase encoding [101], Double-DOE [73], have been proposed along with corresponding reconstruction algorithms over the past few years. In general, great efforts have been made to simplify hyperspectral camera hardware by software reconstruction.
@@ II-B Spectral reconstruction from RGB images
A recent trend to solve the snapshot hyperspectral imaging problem is to exploit hyperspectral data with deep neural networks to reconstruct spectral information from RGB images [4]. Owing to the wide availability of RGB cameras, this approach seems to be a promising candidate for hyperspectral imaging if successful. A large number of neural network architectures have been proposed in the past three NTIRE spectral recovery challenges  [5, 6, 7] and other venues afterwards. Our analysis in this paper focuses on this class of methods to gain insights on their strengths and limitations. In particular, we comprehensively evaluate 17 open-sourced neural networks to date. HSCNN+ [68] is one of the first networks that employs CNN as the backbone for spectral reconstruction which won the first challenge in 2018 [5]. It features dense residual blocks (HCNN-R) and densely-connected structures (HCNN-D). EDSR [55] introduces an enhanced deep super-resolution network by removing unnecessary modules in conventional residual networks. It was originally designed for single-image super-resolution, and later extended for spectral reconstruction. HRNet [105] employs a Hierarchical Regression Network that consists of 4 levels followed by PixelShuffle layers for inter-level interaction, followed by a residual dense block and a residual global block to reconstruct the hyperspectral images. AWAN [52] utilizes a backbone stacked with multiple dual residual attention blocks for dual residual learning. The sensor response function is used as a finer constraint to improve the reconstruction quality. MIRNet [98] adopts a multi-scale residual block that learns contextual information from multiple scales to enhance spatial resolution for image restoration tasks, and later extended for spectral reconstruction. HINet [22] proposes a Half Instance Normalization Block that was originally designed to boost image restoration networks, and later extended for spectral reconstruction. MPRNet [99] is a multi-stage network to learn the projection from degraded measurements to the high-quality images, with a couple of manageable steps. It can also be extended to hyperspectral reconstruction with proper modifications. HDNet [42] proposes a dual domain learning network with a spatial-spectral attention module for pixel-level features, and a frequency domain learning to narrow the frequency domain discrepancy. Restormer [97] was also designed for image restoration tasks initially and then extended for spectral reconstruction, with a focus on designing the building block with the Transformer architecture to capture long-range pixel interactions. MST [14] and MST++ [15] are Transformer-based networks that employ spectral-wise multi-head self-attention to fully make use of spatial sparsity and spectral self-similarity for efficient spectral reconstruction in a coarse-to-fine manner. HySAT [80] employs an exhaustive correlation Transformer to simultaneously model spectral-wise similarity with a token-independent mapping mechanism and particularity with a spectral-wise re-calibration mechanism. HRPN [86] integrates comprehensive multisource priors, in particular the semantic prior of RGB inputs, to regularize and optimize the solution space with a Transformer-based holistic prior-embedded relation network. SSRNet [31] employs a model-guided network based on cross fusion that uses the image formation model and the sensor spectral response function to guide the training of a CNN-backed network. SSTHyper [91] introduces a sparse spectral transformer model to learn shallow and deep spatial-spectral priors and allows adaptive masking of non-significant details. Computational cost is reduced by a cross-level fusion network architecture. MSFN [87] is a multi-stage UNet structure that captures both spatial and spectral features in a multiscale manner. A feature alignment scheme is proposed to preserve spatial correlations and spectral self-similarities. GMSR [83] builds on a more recently developed architecture Mamba [39] as the backbone and develops a lightweight model for global feature representation. Spatial gradient attention and spectral gradient attention are proposed to improve the spectral reconstruction.
II-C Multispectral and hyperspectral image fusion
Another class of related spectral reconstruction methods is to fuse images with low spatial resolution but high spectral resolution with images that have high spatial resolution but low spectral resolution [30, 78]. Different from spectral reconstruction from RGB images, it requires two inputs and a final image with high spatial and spectral resolution is obtained. In recent years, neural networks have also been extensively employed in solving the image fusion problem. Similar datasets as well as remote sensing images are usually used to evaluate the performance. CNN and Transformer are popular backbones in the design of such networks [89, 19, 54, 17, 93]. In this paper, however, we focus on data-driven methods for hyperspectral recovery from single RGB images, where the limitations arising from existing datasets are more prominent.
II-D Dataset bias and data augmentation
@@
II-E Metamerism
Metamerism is a physical phenomenon where distinct spectra produce the same color [2] as the high-dimensional spectral space is projected down to three dimensions of a trichromatic vision system (either the human eye or an RGB camera). This phenomenon has been studied in color science [71, 37], spectral rendering [46, 84, 74], and hyperspectral imaging [40, 36]. In hyperspectral imaging, it is crucial in many applications to distinguish between metamers or near-metamers (i.e., different spectra that project to similar RGB values) [27, 61]. Indeed spectral imaging is usually employed when the RGB color differences between two materials or features are too small to reliably distinguish between them. Therefore, hyperspectral imaging systems require special attention in the system design to acquire accurate spectral signatures [41, 48].
However, modeling metamerism is challenging in data-driven spectral reconstruction as metamerism is hard to capture since they are relatively rare in everyday environments [37], although they are vital for many applications of spectral imaging. Previous works make use of illumination and camera spectral response as means of providing additional information to help improve the spectral reconstruction. For example, Fu et al. [38] propose to select optimal camera response functions from a dataset and recover hyperspectral images with a CNN-based network. This is equivalent to observer metamerism [33] where different cameras observe the spectra differently. Another technique is to explore the structure of illuminant space [2]. For example, Baek et al. [9] employ 29 CIE standard illuminants to augment the hyperspectral dataset for joint hyperspectral and depth reconstruction. Cao et al. [18] develop an unsupervised network to recover spectral information under two lighting conditions. Although such methods enrich the spectral content of hyperspectral datasets, they model metamersim indirectly. Our work differs from these techniques since we directly generate metamers using the metameric black theory [76, 35, 104]. This guarantees the underrepresented metamerism phenomena (the same color from different spectra) in existing datasets can be modeled more effectively.
@@ III-A Spectral image formation
In the past NTIRE challenges [5, 6, 7], optical aberrations have not been included in the image simulation pipeline. However, the optical system of the RGB camera inevitably introduces spectrally-varying blurs to the spectral images, which is modeled as PSFs. This optical process can be described by a linear matrix-vector product in each spectral band followed by a sum over the spectral dimension. The spectral images through the optical system are ${\mathbf{W}}=\mathrm{diag}\left({\mathbf{A}}{\mathbf{X}}\right)$, where ${\mathbf{A}}\in\mathbb{R}^{KMN\times MN}$ is a block matrix that stacks the spectral PSF matrices vertically, and $\mathrm{diag}(\cdot)$ extracts the diagonal blocks. The final RGB image is then
@@ III-B Hyperspectral datasets and data diversity
Compared to very large color (RGB) image datasets (e.g., ImageNet [29], DIV2K [1]), hyperspectral datasets are far smaller in size, primarily limited by the unavailability of high-quality hyperspectral cameras and the difficulty in acquiring outside the lab with moving target scenes. The largest dataset so far is ARAD1K used in the NTIRE 2022 challenge [7]. In addition, we also include the CAVE [96], ICVL [4], and KAUST [53] datasets that share the same spectral range (400 nm to 700 nm) to extend our experiments. The datasets are summarized in Table I. Although other datasets, such as Harvard [21], KAIST [24], and TokyoTech [60] exist, they cover slightly different spectral bands (420 nm to 720 nm), making it difficult to directly compare and cross-validate results among different datasets. We therefore restrict our analysis to the datasets listed in the table.
@@ III-C Modeling metamerism
A spectrum $\mathbf{S}$} can be projected onto two orthogonal subspaces, one for the fundamental metamer $\mathbf{S}$$^{*}$, and the other for metameric black $\mathbf{B}$} [26, 77]. That means the original spectrum can be decomposed as ${\mathbf{S}}={\mathbf{S}}^{*}+{\mathbf{B}}$. The fundamental metamer is a particular solution to Eq. (1), and the metameric black always leads to zero tristimulus, i.e., no impact on the color appearance. The set of all possible metamers is called a metamer set [35]. Wyszecki [88] first introduces a decomposition technique to calculate the metameric black. To generate new metamers, it is possible to add a linear combination of metameric blacks to the fundamental metamer [35]. In a linear algebra perspective, metameric blacks lie in the null space (or kernel) of the camera SRF. Any scalar multiplication to a vector in the null space remains in the null space according to the scalar multiplication property [70]. Inspired by the metameric black theory and the mathematical properties, we propose a simple yet effective way to generate metamers by scaling the metameric black component. A new metameric spectrum $\mathbf{S}$$^{\prime}$ is then
@@
where ${\mathbf{S}}^{*}={\mathbf{Q}}\left({\mathbf{Q}}^{T}{\mathbf{Q}}\right)^{-1}{\mathbf{Q}}^{T}{\mathbf{S}}$ and ${\mathbf{B}}={\mathbf{S}}-{\mathbf{S}}^{*}$. Since adding metameric black does not alter the RGB color, we can vary the coefficient $\alpha$ to generate different spectra that are all metamers. To avoid negative spectral radiance, we clip the negative values in the generated spectra and re-calculate the RGB colors for the affected pixels. See the Appendix for the analysis on the effects of clipping to non-negative values while generating metamer data.
@@ III-D Performance evaluation metrics
${\mathrm{\mathbf{MRAE}}}=\frac{1}{KMN}\sum_{k,i,j}\left|\hat{{\mathbf{X}}}_{k,i,j}-{\mathbf{X}}_{k,i,j}\right|/{\mathbf{X}}_{k,i,j},\\$ (4)
@@ III-E Training details
In our study, we conduct all the experiments across various datasets and network architectures. Following the methodologies proposed by the latest champion network MST++ [15], we employ their patch-wise training approach (patches of $\times128$ pixels).
@@ III-E3 Training and Validation Procedure
[…]
@@ IV Finding 1: Atypical Overfitting
Although it is well-known that deep neural networks may suffer from overfitting problems, we find that the overfitting behavior in spectral reconstruction is atypical and difficult to notice with standard evaluations. Here we introduce minimalist changes to the ARAD1K dataset used in NTIRE 2022 challenge [7] in three experiments to demonstrate it. We exhaustively evaluate a total of 17 open-sourced neural network architectures to-date, namely MST++ [15], MST-L [14], MPRNet [99], Restormer [97], MIRNet [98], HINet [22], HDNet [42], AWAN [52], EDSR [55], HRNet [105], HSCNN+ [68], HySAT [80], HPRN [86], SSTHyper [91], MSFN [87], GMSR [83], and SSRnet [31].
@@ IV-A Training with less data
First, we make a simple change to the training of the participating networks in the NTIRE 2022 challenge [7]. While keeping all the training settings intact, we randomly choose only 50% or 20% of the original training data, respectively, to train the candidate networks and validate the performance on the original validation data. We illustrate the validation curves for MST++ [15] in Fig. 3. See the Supplementary Material for the results of other networks. We summarize the results for 100% and 50% training data in Table II for all networks.
@@
HySAT 100% 0.176 0.028 34.6 0.085
50% 0.254 0.037 31.6 0.089
HPRN 100% 0.257 0.044 30.4 0.098
50% 0.261 0.041 30.8 0.098
SSTHyper 100% 0.181 0.030 33.6 0.083
50% 0.241 0.039 31.1 0.083
MSFN 100% 0.226 0.038 31.9 0.084
50% 0.271 0.044 30.4 0.096
GMSR 100% 0.308 0.056 27.5 0.113
50% 0.393 0.077 25.6 0.127
SSRnet 100% 0.270 0.048 29.5 0.097
50% 0.307 0.052 29.0 0.108
@@ IV-B Validation with unseen data
*CA: chromatic aberration, from a patent double Gauss lens (US20210263286A1).
@@
In our experiments, we generate new RGB images using the same methodology and calibration data, but different noise and compression settings. Specifically, we use the SRF data for a Basler ace 2 camera (model A2a5320-23ucBAS) known to the networks, and simulate Poisson noise at varying noise levels by controlling the number of photon electrons (npe). We adopt the same rudimentary in-camera image signal processing pipeline. As an illustrative example, we use MST++ [15] in Table III, Row 1 as a reference for comparison; results for other networks can be found in the Supplemental Material.
First, as a baseline, we consider a noiseless (npe = 0) and aberration-free case with moderate JPEG compression quality (Q = 65), shown in Table III, Row 2. The results show significant drops in all the performance metrics. Note that the only differences here compared to the challenge dataset are the noise level and compression quality – the base images are identical! This indicates that the network overfits both the noise and JPEG compression parameters.
@@ IV-C Cross-dataset validation
In addition, we inspect the effects of different datasets (cf. Table I) on the performance. We train the MST++ network on the four datasets with the same image simulation parameters. To eliminate the impact of other factors, we choose the ideal noiseless and aberration-free condition without compression. In the validation, we use our trained model on ARAD1K dataset to validate on the other three datasets, respectively. In Table IV, we compare the performance with the models both trained and validated on the original datasets. Results for other networks can be found in the Supplementary Material. They all illustrate the same difficulties in generalization.
TABLE IV: Cross-dataset validation using MST++ [15].
@@
Even though the imaging conditions are the same and ideal, the network trained on one dataset experiences significant performance drops in all metrics when validated on other datasets. This indicates that the contents of the datasets, as well as the acquisition devices used to capture the datasets, play important roles.